The concentration of a biomolecule inside a rod-shaped prokaryotic cell is 0.0017 M. Calculate the number of molecules inside the cell, which is 2.1 ?m long and 0.8 ?m in diameter.
Answer #1

Volume of cell = pi * r^2 h = pi * ( 0.4 * 10^-6)^2 * ( 2.1 * 10^-6) = 1.055 * 10^-18 m^3

Molarity = 0.0017 M

Moles per litre = 0.0017

1 litre = 10^-3 m^3

Moles in 1.055 * 10^-18 m^3 = 1.055 * 10^-18 * 0.0017 / 10^-3 = 1.7944 * 10^-18

Molecules = 6.023 * 10^23 * 1.7944 * 10^-18

= 1.080 * 10^6
Answer #3

Volume, V= pi*r^2*h
    =pi*(0.4*10^-6)^2 * (2.1*10^-6)
    =1.06*10^-18 m^3
    =1.06*10^-18 *(10^6) dm^3
    = 1.06*10^-12 dm^3
   = 1.06*10^-12 L

conc = number of moles/ Volume
0.0017 = number of moles/ 1.06*10^-12
number of moles=1.79*10^-15 moles

number of molecules= number of moles* 6.022*10^23
   = 1.79*10^-15 * 6.022*10^23
   =1.08*10^9 molecules